OpenStudy (anonymous):

WILL GIVE MEDAL Use the ratio test to find the radius of convergence of the power series x + 4 x^2 + 9 x^3 + 16 x^4 + 25 x^5 + \cdots

OpenStudy (anonymous):

what would the equation that you do the ratio test by

OpenStudy (anonymous):

@binarymimic please help

OpenStudy (anonymous):

i know \[a _{n}=C _{n}(x-a)^{n}\]

OpenStudy (anonymous):

where \[(x-a)^{n}=(x)^{n+1}\]

OpenStudy (anonymous):

sorry, was not at computer. can you rewrite this in summation notation? you will find it looks like \[\sum_{k = 1}^{\infty} k ^{2}x ^{k}\] from here take limit of next term divided by previous term.. i.e., \[\lim_{k \rightarrow \infty} \frac{ (k+1)^{2}x ^{k+1} }{ k ^{2}x ^{k} }\]

OpenStudy (anonymous):

should be in absolute value, but it simplifies greatly to just |x|. the ratio test will tell us exactly for what values this series converges

OpenStudy (anonymous):

wouldnt it start at k=0 tho

OpenStudy (anonymous):

i think your first term is just x

OpenStudy (anonymous):

because the first term is x

OpenStudy (anonymous):

k^2 x^k, k = 0 gives 0 as your first term

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

are there like general steps to finding out what the function will look like or do you just have to figure it out differently every time.

OpenStudy (anonymous):

you will have to determine the pattern of the series and write it in a notation such that, given an index, n, you can determine the next term in the sequence (n+1). see http://www.sosmath.com/calculus/radcon/radcon02/radcon02.html i don't know of any general formula, each series is different

OpenStudy (anonymous):

but for this particular one -- the series ratio simplifies to just |x|. that is, as k goes to infinite, the ratio of the next term to the previous term approaches |x|. the ratio test states that, if the ratio of the next term to previous term is < 1, then the ratio converges. in the case where the ratio is |x|, the series converges whenever |x| < 1